Matija Cencelj, Ljubljana, Slovenia

Keywords: smooth free loop space, differentiable slice AMS Subject classification: 58D19

Abstract: Let M be an orientable smooth manifold without boundary. We prove that the differentiable slice theorem for the natural S^{1}-action on the smooth free loop space C^{∞}(S^{1}, M) holds. Poincaré lemma for basic differential forms on C^{∞}(S^{1}, M) is proved.